OpenStudy (anonymous):

if you increase the height by a factor of 3, how much larger will the area be?

OpenStudy (campbell_st):

do a little experiment find the area of a rectangle 2 by 3 multiply 1 side by 3 and leave the other side unchanged what happens to the area

OpenStudy (swissgirl):

let x be the height and y be the length so Area =length*height =x*y=xy So lets say the height is increased by a factor of 3 so our height becomes 3x so area =3x*y=3(xy)

OpenStudy (nincompoop):

laughing out loud length x height ;)

OpenStudy (swissgirl):

hahaha shut up -.- okkkk width *height

OpenStudy (nincompoop):

how about let us ask the girl what shape she thinks she's trying to solve

OpenStudy (campbell_st):

well why do they use base and height for parallelograms and triangles and then length and width for rectangles... makes it all harder than it is @nincompoop

OpenStudy (anonymous):

its a triangle.. and i'm stupid enough y'all have me lost